1. What will be the freezing point of an aqueous solution containing 80.0 g of NaCl in 3.4 L of H2O?
1
Expert's answer
2020-10-20T08:35:42-0400

ΔTf=Kf×Cm

Kf(H2O)=1.86

Cm (molality) is number of moles of solute per kilogram solvent.

ΔTf=Kf×m

M (NaCl) = 58.4 g/mol

n (NaCl) = 80/58.4 = 1.37 mol

Density of water is 1 kg/l. Threfore, mass of 3.4 l of water is 3.4 kg.

Сm (NaCl) = 1.37/3.4=0.4 mol

ΔTf=1.86×0.4=0.74

Theerefore Tf of an aqueous solution containing 80.0 g of NaCl in 3.4 L of H2O will be -0.74°C.